OpenStudy (anonymous):

how would you write this as an equation? 7 units left and flipped upside down

OpenStudy (anonymous):

I don't know what your equation is, so let's use y=x^2 as an example. In order to shift the equation to the left, it'll become y=(x+7)^2. This is because if you were to plug in -7 for your new equation, which would be your new vertex shifted from (0,0), you would get the value of zero. Making a table and checking values will help to ensure that you do the right transformation. In order to flip it upside down, simply add a negative sign to the front of the equation, so your final equation would be y=-(x+7)^2. Hope that helps! As a side note, if you wanted to shift the graph up, you would add the value on the outside, for example, y=(x^2)+3 would shift the graph of x^2 up three units.